# Integer Programming for Generalized Causal Bootstrap Designs

### v2

@@ Abstract
In experimental causal inference, we distinguish between two sources of uncertainty: design uncertainty, due to the treatment assignment mechanism, and sampling uncertainty, when the sample is drawn from a super-population. This distinction matters in settings with small fixed samples and heterogeneous treatment effects, as in geographical experiments. The standard bootstrap procedure most often used by practitioners primarily estimates sampling uncertainty, and the causal bootstrap procedure, which accounts for design uncertainty, was developed for the completely randomized design and the difference-in-means estimator, whereas non-standard designs and estimators are often used in these low-power regimes. We address this gap by proposing an integer program which computes numerically the worst-case copula used as an input to the causal bootstrap method, in a wide range of settings. Specifically, we prove the asymptotic validity of our approach for unconfounded, conditionally unconfounded, and and individualistic with bounded confoundedness assignments, as well as generalizing to any linear-in-treatment and quadratic-in-treatment estimators. We demonstrate the refined confidence intervals achieved through simulations of small geographical experiments.
@@ 1 Introduction
Sampling uncertainty has received extensive attention in the statistics literature. Well-known methods to estimate sampling uncertainty include the bootstrap (Efron,, 1979), in which random draws from the sample mimic the population distribution, and Wald-type confidence intervals, which rely on asymptotic normality of the sample statistic. The methodological toolkit available to quantify design uncertainty is considerably more sparse (Imbens and Menzel,, 2021; Aronow et al.,, 2014). For reasons explained in Appendix 2.1, methods to estimate sampling uncertainty are typically conservative for design uncertainty. The gap widens for experiments with heterogeneous effects, small sample sizes, and a large fraction of the population included in the experiment. Even fewer methods apply to general designs such as those that balance covariates of the treated and control groups, which are widely used precisely in the settings with small experimental samples and heterogeneous treatment effects (Harshaw et al.,, 2024).
In this paper, we propose a novel application of integer programming to identify the joint potential outcome distribution that maximizes the variance of the chosen estimator while being consistent with the randomization design and the observed marginal distributions of potential outcomes. Solving for this joint distribution numerically is what allows us to extend prior work (Aronow et al.,, 2014; Imbens and Menzel,, 2021) which relies on well-known results of the optimality of the isotone copula for the difference-in-means estimator and completely randomized assignment. In particular, our method applies to all unconfounded (and known and probabilistic) assignment mechanisms, as well as any individualistic (and known and probabilistic) assignment mechanisms that verifies some bounded confoundedness. We examine conditionally unconfounded assignments as a special case. Furthermore, our method applies to any linear-in-treatment estimator, i.e. estimators of the form $\sum_{i}Z_{i}a_{i}+b_{i}$, and quadratic-in-treatment estimators, i.e. estimators of the form $\sum_{i}b_{i}+\sum_{j}Z_{i}Z_{j}a_{i,j}$, which include among others, difference-in-means, linear regression on the treatment variable and a covariate, and any doubly robust estimator fit out-of-sample. We provide asymptotic validity results and illustrate the application of our method to a simulated geographical experiment, using real data from the International Monetary Fund (IMF). Our method provides similar coverage and generally tighter confidence intervals across a variety of settings when compared to a widely used, but more conservative baseline.
@@ 1.1 Related works
The quantification of design uncertainty traces its roots to Neyman, (1923), who studied the difference-in-means estimator under complete randomization in the finite population setting. Later work identified tight upper bounds for binary (Robins,, 1988), interval (Stoye,, 2010), and continuous (Aronow et al.,, 2014) outcomes. Imbens and Menzel, (2021) extend the method of Aronow et al., (2014) with a resampling procedure they call the causal bootstrap, providing a more practical confidence interval construction.
@@
Tighter variance estimates can be obtained when predictions are paired with informative covariates. Ding et al., (2019) and Wang et al., (2020) extend the work of Aronow et al., (2014) to the difference-in-means estimator in the presence of covariates, while Fan and Park, (2010) propose the fewer-than-n bootstrap method in the presence of covariates. Another body of literature studies design uncertainty for linear regression estimators. Freedman, (2008), Schochet, (2010), Samii and Aronow, (2012), and Lin, (2013) characterize the contexts in which classical sampling-based variance estimators are valid in the design setting, while Abadie et al., (2014) and Fogarty, (2018) propose regression variance estimators specific to the finite-population setting. Abadie et al., (2020) propose a novel variance estimator for linear regression under both design and sampling uncertainty that trades off between a classical ordinary least squares variance estimator and a novel design-based estimator based on the fraction of the population that is sampled.
@@
Other studies expand design uncertainty estimation beyond Bernoulli or completely randomized designs. Dasgupta et al., (2015) and Lu, (2016) study design uncertainty under factorial designs. Imai, (2008) extends Neyman’s upper bound to the matched-pairs design, which is a special case of covariate-balancing designs; Fogarty, (2018) further extends this result to regression adjustment for matched pairs.
@@
Neyman’s original variance estimator is unbiased only for homogeneous (“strictly additive") treatment effects. Mukerjee et al., (2018) identify a less restrictive set of constraints on the potential outcomes and experimental design under which the variance of certain estimators is unbiasedly estimable. This approach relies on a “Q decomposition” of the variance, where the matrix Q must be derived for each experimental design. The approach fails for certain designs including matched pairs, where units in the same pair cannot be assigned the same treatment. Chattopadhyay and Imbens, (2024) introduce two new variance decompositions for general designs and illustrate their application to completely randomized designs.
@@
Like in (Aronow et al.,, 2014), our approach relies on identifying the variance-maximizing coupling of treated and control observations. We then apply the causal bootstrap method of Imbens and Rubin, (2015) to generate confidence intervals. Our main contribution to these works is to generalize to new estimators and assignment mechanisms through the formulation of an integer program, which no longer requires the worst-case coupling to be known in closed-form. In a related optimization framework, Ji et al., (2023) show how to estimate bounds on a functional of the joint distribution via optimal transport. They do not explicitly solve for the optimal coupling, but their dual approach is robust to misspecification of marginal treated and control outcome probabilities in the presence of covariates. Finally, Harshaw et al., (2021) propose a minimal variance bound among all quadratic forms through a convex optimization procedure. Their method yields interesting admissibility results but only considers quadratic bounds. Our approach, inspired by the Frechet-Hoeffding bound approach used by Aronow et al., (2014) and Imbens and Menzel, (2021), avoids this limitation.
@@ 2.1 Notation and background
To build intuition, we briefly review Neyman’s conservative bound and follow-up work on the difference-in-means estimator under a completely randomized assignment. We consider a finite-population potential outcomes model in which each unit $i\in[N]$ is assigned to treatment ($Z_{i}=0$) or control ($Z_{i}=1$), and outcome $Y_{i}(Z_{i})$ is observed. The average treatment effect is defined as $\tau=\tfrac{1}{N}\sum_{i\in[N]}Y_{i}(1)-Y_{i}(0)$. Let $N_{1}$ and $N_{0}$ be the number of treated and controlled units respectively. Define the difference in means estimator $\hat{\tau}_{\textrm{DIM}}=\tfrac{1}{N_{1}}\sum_{i:Z_{i}=1}Y_{i}-\tfrac{1}{N_{0}}\sum_{i:Z_{i}=0}Y_{i}$ and the completely randomized design that assigns exactly $N_{1}$ units to treatment and $N_{0}$ units to control. Neyman, (1923) decomposed the variance of this estimator as
@@
where $S_{z}^{2}$ is the sample variances of the treated ($z=1$) and control ($z=0$) units, while $S_{\tau}^{2}$ is the sample variance of the unit-level treatment effects $Y_{i}(1)-Y_{i}(0)$. See Appendix 7.9 for the full formulas. $S_{\tau}^{2}$ depends on the unknown joint distribution between treated and control outcomes since it contains terms $Y_{i}(1)Y_{i}(0)$, making it impossible to estimate consistently from experimental data in which only one of $Y_{i}(1)$ or $Y_{i}(0)$ is ever observed. While Neyman used the fact that variances are nonnegative to conservatively bound $S_{\tau}^{2}\geq 0$, thus proposing an upper-bound of $\mathbf{Var}_{Z}[\hat{\tau}_{\textrm{DIM}}]$, this bound can be improved by identifying a joint distribution that maximizes $S_{\tau}^{2}$ while matching the observed marginal distributions of $Y(0)$ and $Y(1)$. Aronow et al., (2014) identify this variance-maximizing distribution to be the isotone (assortative) copula, which pairs the smallest treated outcome with the smallest control outcome, then the second-smallest, and so on. They derive analytical variance bounds using that joint distribution, while Imbens and Menzel, (2021) propose a bootstrap-style approach, imputing the counterfactual outcomes using the isotone copula and resampling those imputed outcomes to generate bootstrap samples.
@@ 2.2 Our approach
There are multiple ways of constructing confidence intervals from the solution to this optimization objective. We propose to follow a similar procedure to the causal bootstrap method proposed by Imbens and Menzel, (2021): we impute unobserved potential outcomes using the least-favorable copula obtained above and drawing bootstrap samples via rerandomization of units to treatment. Confidence intervals are then constructed from the quantiles of this causal bootstrap distribution. We refer the reader to a full discussion in Appendix 7.10.
@@
where $\mathbf{X}\in\{0,1\}^{(2\cdot N\cdot K)}$ and $\mathbf{Y}\in\mathbb{R}^{N\times(2N\cdot K)}$ are a vector and matrix representation of the indicator variables and potential outcomes respectively; cf. Appendix 7.3 for details.
@@ 2.3 Adding constraints on the potential outcomes
A clever observation by Aronow et al., (2014) is that the observed potential outcomes allow us to impose further constraints on the missing potential outcomes. While they focus entirely on the completely randomized assignment, which we use now to build intuition, we will generalize our results to other assignments in Section 4.1. If all units receive treatment with the same probability (different from 1 or 0), we expect the treated observed $F^{obs}_{1}$ and missing $F^{mis}_{1}$ marginals to match and the controlled observed $F^{obs}_{0}$ and missing $F^{mis}_{0}$ marginals to match. Formally, for any outcome $y$ and treatment assignment value $a\in\{0,1\}$, let $F^{obs}_{a}(y):=\frac{1}{N_{a}}\sum_{i=1}^{N}\mathbb{I}(Z_{i}=a,Y_{i}(a)=y)$ and $F^{mis}_{a}(y):=\frac{1}{N_{1-a}}\sum_{i=1}^{N}\mathbb{I}(Z_{i}=1-a,Y_{i}(a)=y)$ For a completely randomized assignment, each observed and missing marginal is equal in expectation to the true underlying marginal $F_{a}$. In other words, we have the following equality for any potential outcome $y$:
@@ 2.4 Our proposed integer linear program
When indicated, the equations above are defined for all $a\in\{0,1\}$, for all $i\in[1,N]$, and for all $y_{k}\in\mathcal{Y}$. As written, the optimization has binary variables and its constraints are all linear, but the objective is quadratic. Because the product of two binary variables is also a binary variable, the objective can be rewritten as a linear function of binary variables, for which powerful solvers are available, using standard transformations; details are provided in Appendix 7.3.
@@
The regime of feasibility is captured in the following Lemma, a proof of which can be found in Appendix 7.2.
Lemma 1.
@@ 3.1 Linear-in-treatment estimators
So far, our results hold only for the difference-in-means estimator, which has been the primary object of study of previous work. In fact, they can be easily extended to a slightly more general class of estimators of the form: $\hat{\tau}:=\sum_{i=1}^{N}Z_{i}a_{i}+b_{i}$ , where $a_{i}$ and $b_{i}$ are constants with respect to the assignment mechanism and linear in the potential outcomes. We refer to such estimators as being linear-in-treatment (the second linearity assumption is then implied, for simplicity). Whether an estimator is linear-in-treatment depends on the assignment mechanism. Strictly speaking, the difference-in-means estimator is linear-in-treatment for a completely randomized assignment but not for a Bernoulli assignment, since the normalization constants $N_{1}^{-1}$ and $N_{0}^{-1}$ are not constants with respect to $Z$. The Horvitz-Thompson estimator (Horvitz and Thompson,, 1952) is linear-in-treatment for either assignment mechanism:
@@ 3.2 Fitting regressions out-of-sample
There are many practical settings in which fitted-out-of-sample functions are available. For example, such functions may be fitted on the same cohort before the start of the experiment. A typical and straightforward example is to use a pre-treatment-period value of the outcome as a plug-in estimator (Abadie,, 2005; Deng et al.,, 2013). If we expect growth of this variable, we can also account for its rate of growth and rescale the plug-in estimator with its expected value; however, fitting this growth rate using data from the experimental period depends on the treatment assignment $\mathbf{Z}$ and should be avoided. We include an example of this approach in Section 5. Another option is to rely on sample-splitting, where we split units into $C$ cohorts, chosen independently of the treatment assignment $\mathbf{Z}$. In that case, we can estimate a (pair of) function(s) per cohort using data from all other cohorts.

### v3

@@ Abstract
In experimental causal inference, we distinguish between two sources of uncertainty: design uncertainty, due to the treatment assignment mechanism, and sampling uncertainty, when the sample is drawn from a super-population. This distinction matters in settings with small fixed samples and heterogeneous treatment effects, as in geographical experiments. The standard bootstrap procedure most often used by practitioners primarily estimates sampling uncertainty, and the causal bootstrap procedure, which accounts for design uncertainty, was developed for the completely randomized design and the difference-in-means estimator, whereas non-standard designs and estimators are often used in these low-power regimes. We address this gap by proposing an integer program which computes numerically the worst-case copula used as an input to the causal bootstrap method, in a wide range of settings. Specifically, we prove the asymptotic validity of our approach for unconfounded, conditionally unconfounded, and individualistic with bounded confoundedness assignments, as well as generalizing to any linear-in-treatment and quadratic-in-treatment estimators. We demonstrate the refined confidence intervals achieved through simulations of small geographical experiments.
causal inference, experiment design, bootstrap
@@ 1 Introduction
Sampling uncertainty has received extensive attention in the statistics literature. Well-known methods to estimate sampling uncertainty include the bootstrap (Efron, 1979), in which random draws from the sample mimic the population distribution, and Wald-type confidence intervals, which rely on asymptotic normality of the sample statistic. The methodological toolkit available to quantify design uncertainty is considerably more sparse (Imbens & Menzel, 2021; Aronow et al., 2014). For reasons explained in Appendix 2.1, methods to estimate sampling uncertainty are typically conservative for design uncertainty. The gap widens for experiments with heterogeneous effects, small sample sizes, and a large fraction of the population included in the experiment. Even fewer methods apply to general designs such as those that balance covariates of the treated and control groups, which are widely used precisely in the settings with small experimental samples and heterogeneous treatment effects (Harshaw et al., 2024).
In this paper, we propose a novel application of integer programming to identify the joint potential outcome distribution that maximizes the variance of the chosen estimator while being consistent with the randomization design and the observed marginal distributions of potential outcomes. Solving for this joint distribution numerically is what allows us to extend prior work (Aronow et al., 2014; Imbens & Menzel, 2021) which relies on well-known results of the optimality of the isotone copula for the difference-in-means estimator and completely randomized assignment. In particular, our method applies to all unconfounded (and known and probabilistic) assignment mechanisms, as well as any individualistic (and known and probabilistic) assignment mechanisms that verifies some bounded confoundedness. We examine conditionally unconfounded assignments as a special case. Furthermore, our method applies to any linear-in-treatment estimator, i.e. estimators of the form $\sum_{i}Z_{i}a_{i}+b_{i}$, and quadratic-in-treatment estimators, i.e. estimators of the form $\sum_{i}b_{i}+\sum_{j}Z_{i}Z_{j}a_{i,j}$, which include among others, difference-in-means, linear regression on the treatment variable and a covariate, and any doubly robust estimator fit out-of-sample. We provide asymptotic validity results and illustrate the application of our method to a simulated geographical experiment, using real data from the International Monetary Fund (IMF). Our method provides similar coverage and generally tighter confidence intervals across a variety of settings when compared to a widely used, but more conservative baseline.
@@ 1.1 Related works
The quantification of design uncertainty traces its roots to Neyman (1923), who studied the difference-in-means estimator under complete randomization in the finite population setting. Later work identified tight upper bounds for binary (Robins, 1988), interval (Stoye, 2010), and continuous (Aronow et al., 2014) outcomes. Imbens & Menzel (2021) extend the method of Aronow et al. (2014) with a resampling procedure they call the causal bootstrap, providing a more practical confidence interval construction.
@@
Tighter variance estimates can be obtained when predictions are paired with informative covariates. Ding et al. (2019) and Wang et al. (2020) extend the work of Aronow et al. (2014) to the difference-in-means estimator in the presence of covariates, while Fan & Park (2010) propose the fewer-than-n bootstrap method in the presence of covariates. Another body of literature studies design uncertainty for linear regression estimators. Freedman (2008), Schochet (2010), Samii & Aronow (2012), and Lin (2013) characterize the contexts in which classical sampling-based variance estimators are valid in the design setting, while Abadie et al. (2014) and Fogarty (2018) propose regression variance estimators specific to the finite-population setting. Abadie et al. (2020) propose a novel variance estimator for linear regression under both design and sampling uncertainty that trades off between a classical ordinary least squares variance estimator and a novel design-based estimator based on the fraction of the population that is sampled.
@@
Other studies expand design uncertainty estimation beyond Bernoulli or completely randomized designs. Dasgupta et al. (2015) and Lu (2016) study design uncertainty under factorial designs. Imai (2008) extends Neyman’s upper bound to the matched-pairs design, which is a special case of covariate-balancing designs; Fogarty (2018) further extends this result to regression adjustment for matched pairs.
@@
Neyman’s original variance estimator is unbiased only for homogeneous (“strictly additive”) treatment effects. Mukerjee et al. (2018) identify a less restrictive set of constraints on the potential outcomes and experimental design under which the variance of certain estimators is unbiasedly estimable. This approach relies on a “Q decomposition” of the variance, where the matrix Q must be derived for each experimental design. The approach fails for certain designs including matched pairs, where units in the same pair cannot be assigned the same treatment. Chattopadhyay & Imbens (2024) introduce two new variance decompositions for general designs and illustrate their application to completely randomized designs.
@@
Like in (Aronow et al., 2014), our approach relies on identifying the variance-maximizing coupling of treated and control observations. We then apply the causal bootstrap method of Imbens & Rubin (2015) to generate confidence intervals. Our main contribution to these works is to generalize to new estimators and assignment mechanisms through the formulation of an integer program, which no longer requires the worst-case coupling to be known in closed-form. In a related optimization framework, Ji et al. (2023) show how to estimate bounds on a functional of the joint distribution via optimal transport. They do not explicitly solve for the optimal coupling, but their dual approach is robust to misspecification of marginal treated and control outcome probabilities in the presence of covariates. Finally, Harshaw et al. (2021) propose a minimal variance bound among all quadratic forms through a convex optimization procedure. Their method yields interesting admissibility results but only considers quadratic bounds. Our approach, inspired by the Frechet-Hoeffding bound approach used by Aronow et al. (2014) and Imbens & Menzel (2021), avoids this limitation.
@@ 2.1 Notation and background
To build intuition, we briefly review Neyman’s conservative bound and follow-up work on the difference-in-means estimator under a completely randomized assignment. We consider a finite-population potential outcomes model in which each unit $i\in[N]$ is assigned to treatment ($Z_{i}=0$) or control ($Z_{i}=1$), and outcome $Y_{i}(Z_{i})$ is observed. The average treatment effect is defined as $\tau=\tfrac{1}{N}\sum_{i\in[N]}Y_{i}(1)-Y_{i}(0)$. Let $N_{1}$ and $N_{0}$ be the number of treated and controlled units respectively. Define the difference in means estimator $\hat{\tau}_{\textrm{DIM}}=\tfrac{1}{N_{1}}\sum_{i:Z_{i}=1}Y_{i}-\tfrac{1}{N_{0}}\sum_{i:Z_{i}=0}Y_{i}$
and the completely randomized design that assigns exactly $N_{1}$ units to treatment and $N_{0}$ units to control. Neyman (1923) decomposed the variance of this estimator as
@@
where $S_{z}^{2}$ is the sample variances of the treated ($z=1$) and control ($z=0$) units, while $S_{\tau}^{2}$ is the sample variance of the unit-level treatment effects $Y_{i}(1)-Y_{i}(0)$. See Appendix A.8 for the full formulas. $S_{\tau}^{2}$ depends on the unknown joint distribution between treated and control outcomes since it contains terms $Y_{i}(1)Y_{i}(0)$, making it impossible to estimate consistently from experimental data in which only one of $Y_{i}(1)$ or $Y_{i}(0)$ is ever observed. While Neyman used the fact that variances are nonnegative to conservatively bound $S_{\tau}^{2}\geq 0$, thus proposing an upper-bound of $\mathbf{Var}_{Z}[\hat{\tau}_{\textrm{DIM}}]$, this bound can be improved by identifying a joint distribution that maximizes $S_{\tau}^{2}$ while matching the observed marginal distributions of $Y(0)$ and $Y(1)$. Aronow et al. (2014) identify this variance-maximizing distribution to be the isotone (assortative) copula, which pairs the smallest treated outcome with the smallest control outcome, then the second-smallest, and so on. They derive analytical variance bounds using that joint distribution, while Imbens & Menzel (2021) propose a bootstrap-style approach, imputing the counterfactual outcomes using the isotone copula and resampling those imputed outcomes to generate bootstrap samples.
@@ 2.2 Our approach
There are multiple ways of constructing confidence intervals from the solution to this optimization objective. We propose to follow a similar procedure to the causal bootstrap method proposed by Imbens & Menzel (2021): we impute unobserved potential outcomes using the least-favorable copula obtained above and drawing bootstrap samples via rerandomization of units to treatment. Confidence intervals are then constructed from the quantiles of this causal bootstrap distribution. We refer the reader to a full discussion in Appendix A.9.
@@
where $\mathbf{X}\in\{0,1\}^{(2\cdot N\cdot K)}$ and $\mathbf{Y}\in\mathbb{R}^{N\times(2N\cdot K)}$ are a vector and matrix representation of the indicator variables and potential outcomes respectively; cf. Appendix A.3 for details.
@@ 2.3 Adding constraints on the potential outcomes
A clever observation by Aronow et al. (2014) is that the observed potential outcomes allow us to impose further constraints on the missing potential outcomes. While they focus entirely on the completely randomized assignment, which we use now to build intuition, we will generalize our results to other assignments in Section 4.1. If all units receive treatment with the same probability (different from 1 or 0), we expect the treated observed $F^{obs}_{1}$ and missing $F^{mis}_{1}$ marginals to match and the controlled observed $F^{obs}_{0}$ and missing $F^{mis}_{0}$ marginals to match. Formally, for any outcome $y$ and treatment assignment value $a\in\{0,1\}$, consider
$F^{obs}_{a}(y)$ $:=\frac{1}{N_{a}}\sum_{i=1}^{N}\mathbb{I}(Z_{i}=a,Y_{i}(a)=y)\,,$
$F^{mis}_{a}(y)$ $:=\frac{1}{N_{1-a}}\sum_{i=1}^{N}\mathbb{I}(Z_{i}=1-a,Y_{i}(a)=y)\,.$
For a completely randomized assignment, each observed and missing marginal is equal in expectation to the true underlying marginal $F_{a}$. In other words, we have the following equality for any potential outcome $y$:
@@ 2.4 Our proposed integer linear program
When indicated, the equations below are defined for all $a\in\{0,1\}$, for all $i\in[1,N]$, and for all $y_{k}\in\mathcal{Y}$. As written, the optimization has binary variables and its constraints are all linear, but the objective is quadratic. Because the product of two binary variables is also a binary variable, the objective can be rewritten as a linear function of binary variables, for which powerful solvers are available, using standard transformations; details are provided in Appendix A.3.
@@
The regime of feasibility is captured in the following Lemma, a proof of which can be found in Appendix A.2.
Lemma 2.1.
@@ 3.1 Linear-in-treatment estimators
So far, our results hold only for the difference-in-means estimator, which has been the primary object of study of previous work. In fact, they can be easily extended to a slightly more general class of estimators of the form: $\hat{\tau}:=\sum_{i=1}^{N}Z_{i}a_{i}+b_{i}$ , where $a_{i}$ and $b_{i}$ are constants with respect to the assignment mechanism and linear in the potential outcomes. We refer to such estimators as being linear-in-treatment (the second linearity assumption is then implied, for simplicity). Whether an estimator is linear-in-treatment depends on the assignment mechanism. Strictly speaking, the difference-in-means estimator is linear-in-treatment for a completely randomized assignment but not for a Bernoulli assignment, since the normalization constants $N_{1}^{-1}$ and $N_{0}^{-1}$ are not constants with respect to $Z$. The Horvitz-Thompson estimator (Horvitz & Thompson, 1952) is linear-in-treatment for either assignment mechanism:
@@ 3.2 Fitting regressions out-of-sample
There are many practical settings in which fitted-out-of-sample functions are available. For example, such functions may be fitted on the same cohort before the start of the experiment. A typical and straightforward example is to use a pre-treatment-period value of the outcome as a plug-in estimator (Abadie, 2005; Deng et al., 2013). If we expect growth of this variable, we can also account for its rate of growth and rescale the plug-in estimator with its expected value; however, fitting this growth rate using data from the experimental period depends on the treatment assignment $\mathbf{Z}$ and should be avoided. We include an example of this approach in Section 5. Another option is to rely on sample-splitting, where we split units into $C$ cohorts, chosen independently of the treatment assignment $\mathbf{Z}$. In that case, we can estimate a (pair of) function(s) per cohort using data from all other cohorts.